OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -2) and (-1,0)?

OpenStudy (anonymous):

m=(0-(-2))/(-3-3)=(2)/(-6)= -1/3

OpenStudy (anonymous):

Remember slope = (difference in y)/(difference in x).

OpenStudy (anonymous):

or rise/run was always taught to me

OpenStudy (anonymous):

\[(y - y )/ (x - x)\]

OpenStudy (anonymous):

What is the y in the first point?

OpenStudy (anonymous):

@Slowly_Changing Is the y 3 or -2?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

That's right, now what is the y in the second point?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Good, so now you have the first part. (-2 - 0) Now what is the x in the first point? What is the x in the second point?

OpenStudy (anonymous):

When you have the two y's or x's, it does not matter which order you subtract them. You could do -2 - 0 like I did, or you could do 0 - (-2). It doesn't matter.

OpenStudy (anonymous):

@Slowly_Changing

OpenStudy (anonymous):

What are the two x's?